OpenStudy (anonymous):

How would I solve this: Solve for x: 3(x + 1) = −2(x − 1) − 4

OpenStudy (anonymous):

open the parentheses: \[3x + 3 = -2x + 2 -4\] Did you understand how i did it?

OpenStudy (anonymous):

You distributed out the numbers

OpenStudy (anonymous):

Do you want me to continue or you can take it from here?

OpenStudy (anonymous):

Continue

OpenStudy (anonymous):

\[3x + 3 = -2x + 2 -4\] now: \[3x + 2x = 2 - 4 -3\] and then: \[5x = -5\] so the answer is: \[x = -1\]

OpenStudy (anonymous):

thank you so much I understand now

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

Wait How did you get the second answer after 3x+3=−2x+2−4

OpenStudy (anonymous):

I moved all the variables to the left and isolated them.